How many real solutions does this equation have?
(x^2-9x+19)^(2x^3-x^2-10x)=1

Answer by stanbon(75887)

(x^2-9x+19)^(2x^3-x^2-10x)=1
For this to be true the exponent must be 0
so 2x^3-x^2-10x=0
x(2x^2-x-10)=0
x(2x-5)(x+2)=0
x=0 or x=5/2 or x=-2
Cheers
Stan H.